# Hybrid inflation and gravitational waves from accidentally light scalars

### v1

@@ 5 Topological defects
Magnetic monopoles can appear if the second homotopy group of the vacuum manifold is nonzero. When $G=\mathrm{SO}(3)$ is broken to $\mathrm{SO}(2)$ by a triplet VEV, as is the case in our model, this gives rise to the celebrated ’t Hooft-Polyakov monopole [39, 40] with $\pi_{2}(\mathrm{SO}(3)/\mathrm{SO}(2))=\pi_{2}[S^{2}]=\mathbb{Z}$. However, in our models the relevant phase transition takes place before inflation, and therefore no monopoles are likely to remain in our Hubble patch. To be precise, symmetry breaking in our model proceeds in two steps: first, at the scale $v$, the symmetry group $G$ is broken completely by the 5-plet VEV, while the triplet VEV is zero in this phase. The vacuum manifold (or rather the scalar field manifold which minimizes the potential for all fields except the slowly-rolling inflaton) is $G$, and in different causally connected regions of the universe, the scalar fields will take VEVs in different domains of $G$. However, the subsequent phase of inflation stretches out these inhomogeneities, such that, at all scales relevant to us, the $5$-plet VEV becomes effectively a single point in field space. Nontrivial field configurations which might evolve into monopoles are “inflated away”. The second step of symmetry breaking happens at the waterfall transition at the end of inflation, when the $5$-plet VEV relaxes to zero and a triplet VEV is generated, thereby restoring an $\mathrm{SO}(2)$ symmetry. The vacuum manifold becomes $\mathrm{SO}(3)/\mathrm{SO}(2)\simeq S^{2}$; however, no new monopoles can be created, since this would require energies sufficiently high to restore the entire $\mathrm{SO}(3)$ symmetry. We conclude that, for better or worse, magnetic monopoles will not affect the cosmology of our model.
@@
Cosmic strings can appear if the scalar vacuum manifold is not simply connected. This is not the case in the minimal model, but it is interesting to consider a non-minimal extension to study the phenomenological consequences. Consider a version of the model in which the triplet $\chi$ is a complex scalar field transforming under $G=\mathrm{SU}(2)\times\mathrm{U}(1)_{\chi}$, with $\mathrm{U}(1)_{\chi}$ acting as a phase rotation. All or part of $G$ may be thought of as gauged. We can further extend the model by promoting also $\phi$ to a complex field and gauging $\mathrm{U}(1)_{\phi}$, which is appealing because then the model no longer relies on any ad-hoc discrete symmetries, the role of the $\mathbb{Z}_{2}$ in the original model now being played by $\mathrm{U}(1)_{\phi}$. The technical details of these model versions are worked out in Appendix A. These models do allow for additional quartic scalar couplings, but the qualitative features of the inflaton potential, as well as CMB predictions, do not change: crucially, there is still an accidentally light scalar field which can play the role of the inflaton, with a cosinusoidal loop-induced potential analogous to Eq. (6). Likewise, the effective mass parameter of the waterfall field $\chi$ depends on the value of the inflaton, and parameters can be chosen such that it becomes tachyonic at the end of inflation. The waterfall field is zero along the inflationary trajectory, so $\mathrm{U}(1)_{\chi}$ remains unbroken during inflation while $\mathrm{SU}(2)\times\mathrm{U}(1)_{\phi}$ is completely broken. The true vacuum is at $\langle\phi\rangle=0$ and $\langle|\chi|\rangle=\langle|\chi^{3}|\rangle=v_{\chi}\neq 0$, so the vacuum manifold is $S^{2}\times\mathrm{U}(1)_{\chi}$.
@@
where $v_{\chi}$ is the breaking scale of $\mathrm{U}(1)_{\chi}$. It is well known that this network quickly reaches a scaling regime, where the fraction of energy density stored in the CSs redshifts in the same way as the energy density of the background [42, 41]. Strings can intercommute and produce loops, which oscillate under their own tension and radiate energy in the form of GWs. Such a signal of GWs appears to us as a SGWB with a characteristic scale-invariant spectrum (if standard cosmology is assumed [43]). The peak and the amplitude of the GWs signal for stable strings are both given by the string tension, which, in order to match the signal recently reported by the NANOGrav collaboration, has to be $G_{\mathrm{\scriptscriptstyle{N}}}\,\mu\simeq 10^{-10}$, even though stable CSs seem to be disfavoured with respect to other cosmological sources [11]. The reason is that, for stable CSs, both the amplitude and the shape of the GW signal are controlled by $G_{\mathrm{\scriptscriptstyle{N}}}\mu$; once we fix the string tension to match the amplitude of the reported signal, the CS spectrum turns out to be flat in the nHz region, with data preferring a blue-tilted spectrum. $G_{\mathrm{\scriptscriptstyle{N}}}\mu\simeq 10^{-10}$ represents a trade-off between a signal which is not too weak and one which is not too flat. However, as the computation of the expected signal of GWs from a CS network is subject to large uncertainties, we can take as upper bound on the string tension the one given by the CMB constraint on the impact of CSs on the CMB power spectrum [44], $G_{\mathrm{\scriptscriptstyle{N}}}\mu\lesssim 10^{-7}$, which corresponds to an inflationary scale $V_{0}^{1/4}\lesssim\left(\lambda_{\chi}+\lambda_{\chi}^{\prime}\right)\left(3\times 10^{14}\right)\,{\rm GeV}$.
Finally, domain walls can appear if the scalar vacuum manifold is not connected. This, again, is not the case in our minimal model. In the following we exploit a non-minimal model with disconnected minima, which is described in detail in Appendix B. In this case the PTA signal can be fit with GWs produced from annihilating DWs, as explained in the next section. The waterfall field $\chi$ is taken to be a complex scalar triplet, subject to a discrete symmetry $\mathbb{Z}_{4\chi}:\chi\,\rightarrow\,i\chi$. This leads to disconnected minima which are degenerate, corresponding to stable DWs. However, the vacuum degeneracy can be easily lifted by softly breaking the $Z_{4\chi}$ discrete symmetry, allowing for the DWs to annihilate. Also in this model the inflationary predictions are the same as those for the minimal model.
@@ 5.1 Gravitational waves from unstable domain walls
In the model described in Appendix B, after the end of inflation the $\chi$ scalar potential has two disconnected, degenerate minima, so that DWs are formed with a surface energy density [41]
@@
However, the discrete symmetry ensuring the vacuum degeneracy may be softly broken, which lifts the degeneracy and introduces a potential bias $\Delta V$, see Eq. (B6). Provided that $\Delta V<0.795\,V_{0}$ [45], where $V_{0}$ is the potential energy at $\chi=0$, DWs are formed, but they are unstable as long as $\Delta V\neq 0$. The different domains annihilate when the bias pressure $\Delta V$ dominates over the DW surface tension at
@@
with $C_{d}\simeq 3$ [47], until all the observable Universe lies in the true vacuum. We require the DW network to disappear before the bias energy density of the false-vacuum patches dominates the Universe,
@@
Here $R\left(t_{\rm ann}\right)$ and $R\left(t_{0}\right)$ are the scale factors at the time of DW annihilation and today, respectively, $g_{*}$ is the number of effective degrees of freedom in relativistic energy, and $T_{\rm ann}$ is the temperature of the Universe at the moment of DW annihilation,
@@
Numerical simulations have shown that $b,c\simeq 1$ for $\mathbb{Z}_{2}$-type DWs [46], and our model falls in this category, while causality fixes $a=3$ [48]. [^7] In principle, the causality tail of the GW spectrum is affected by the QCD phase transition and the subsequent change in the number of relativistic degrees of freedom. This leads to a deviation from the pure power-law behaviour with $a=3$. Including the effect of the QCD phase transition is beyond the scope of this work and we refer to Ref. [49] for a detailed discussion on the topic. By plugging Eq. (35) and Eq. (38) into Eq. (37), we observe that the GW peak frequency scales with the DW network parameters as $f_{p}\sim\sqrt{\Delta V/\sigma}$. Similarly, by plugging Eq. (40) into Eq. (39), we find that $h^{2}\Omega_{\rm GW}(f_{p})\sim(\sigma^{2}/\Delta V)^{2}$.
By taking the quartic couplings to be of order one, the DW tension is related to the inflationary scale as $V_{0}\simeq\sigma^{4/3}$. This allows us to relate the GW signal generated by the DW network to inflation. It has been shown [11, 47, 50] that a network of unstable DWs is a good candidate to explain the signal of a SGWB recently detected by the NANOGrav collaboration [11]. We find that a model with $V_{0}^{1/4}\simeq 2\times 10^{5}\,{\rm GeV}$ and a suitable bias $\Delta V$ leads to a GW signal which fits the NANOGrav detection. The peak of the GW signal is within the reach of the upcoming space-borne interferometer LISA [51] for $3\times 10^{7}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 2\times 10^{10}\,{\rm GeV}$, while for $3\times 10^{10}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 7\times 10^{12}\,{\rm GeV}$ the SGWB might be detected by the ET in the future. Fig. 5 shows several different benchmark GW spectra together with the sensitivity curves of current and planned experiments.
The energy density released by DW annihilation can potentially alter the expansion rate of the Universe. In order not to spoil BBN, the products of DW annihilation have to decay before the BBN epoque, either into dark radiation (DR) or into SM particles [52]. Indeed, $\chi$ can decay into the $\mathrm{SO}(2)$ dark photon, or into SM degrees of freedom via e.g. a Higgs-portal coupling. The dominant decay channel depends on the relative magnitude of the couplings and is related to the reheating mechanism, which we do not explore in this paper. The usual assumption, which also leads to the weakest constraints, is that such decays are already active at $T_{\rm ann}$ [52]. Assuming the DW energy density is transferred entirely into DR, the extra number of neutrinos species receives a correction [53]
@@ 6 Conclusions
Models with accidentally light scalars [3] are particularly well suited for realising hybrid inflation. They feature tree-level flat directions, lifted by loop corrections, along which slow-roll inflation takes place. They also feature tree-level couplings of the inflaton to other scalar fields, which can become tachyonic along the inflationary direction, thereby ending inflation by the waterfall mechanism.
In this paper we have presented a simple yet realistic example. It is defined by a symmetry group $G=\mathrm{SO}(3)$ and two scalar multiplets $\phi$ and $\chi$, transforming in the representation $\boldsymbol{5}$ and $\boldsymbol{3}$ respectively. The two fields may also carry an abelian charge (discrete or continuous), depending on the variant of the model. Gauging $\mathrm{SO}(3)$ and allowing for all renormalizable potential terms compatible with the symmetry, one degree of freedom $a$ in $\phi$ remains massless at tree level: this is what we call an accidentally light field, which we identify with the inflaton. The inflationary potential arises at the one-loop level, induced by the gauge coupling and a quartic coupling between $\phi$ and $\chi$. It has a form similar to that of Natural Inflation [4]. One component of $\chi$ plays the role of the waterfall field: its mass-squared starts out positive and decreases during inflation until, at some critical point along the inflationary direction, it turns tachyonic. This destabilizes the waterfall field which terminates inflation by fast-rolling down its steep potential.
The flatness of the inflationary potential is naturally protected: it is guaranteed by symmetry at the tree level and it gets lifted by one-loop radiative corrections. Unlike pseudo-Goldstone inflatons, here the inflaton has tree-level non-derivative couplings to other scalars, which are necessary for an efficient waterfall. These couplings are typically a concern in earlier realisations of hybrid natural inflation, as they spoil the flatness, unless they are forbidden by some complicated additional symmetries or fine-tuned.
@@
Inflation ends as the waterfall field becomes tachyonic, and the Universe is reheated with a reheating temperature of the order of the inflationary scale, $T_{\rm RH}\sim{\cal O}(V_{0}^{1/4})$. A negative curvature of the potential triggers tachyonic instability of the waterfall field quantum fluctuations, which are exponentially enhanced. This leads to production of GWs, which reach us as a SGWB signal. The peak frequency of the GW spectrum is typically as large as $\mathcal{O}\left(10^{10}\right)\,{\rm Hz}$ but it can be lowered by sufficiently decreasing the inflationary scale. We find that the signal of GWs produced during tachyonic instability is within the reach of the Einstein Telescope for an inflationary scale in the range $3\times 10^{5}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 5\times 10^{11}{\rm GeV}$. Such a low inflationary scale, despite pointing to some UV completion well below $M_{\mathrm{P}}$, is perfectly viable and is associated to an unobservably small tensor-to-scalar ratio. This confirms that GWs from a tachyonic instability can act as a possible probe of low-scale inflationary scenarios.
The inflationary phase may be followed by a production of topological defects, depending on the specific symmetries which undergo SSB. In a model variation with $\phi$ and $\chi$ charged under an additional $U(1)$ gauge symmetry, the vacuum manifold is such that CSs are produced during the waterfall transition. The tension of the string network is bounded by CMB, and this constrains the inflationary scale to be $V_{0}^{1/4}\lesssim$ a few ×$\times 10^{14}\,{\rm GeV}$. Another model variation features a softly-broken $\mathbb{Z}_{4}$ symmetry acting on $\phi$ and $\chi$. In this case the vacuum manifold is disconnected, and unstable DWs are generated at the end of inflation. They produce a SGWB, which can fit the signal recently detected by the NANOGrav collaboration [11], albeit for an extremely low inflationary scale, $V_{0}^{1/4}\simeq 2\times 10^{5}\,{\rm GeV}$, assuming order one quartic couplings. The DW annihilation signal can also be relevant for future detection by LISA (ET) for $3\times 10^{7}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 2\times 10^{10}\,{\rm GeV}$ ($3\times 10^{10}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 7\times 10^{12}\,{\rm GeV}$).
Upcoming experiments will soon be able to further constrain or rule out our models, for example, by measuring the sign of $n_{r}$, or by a discovery of a SGWB by the LIGO-Virgo-KAGRA network. Concerning future directions, it would be interesting to couple our models to the Standard Model in order to study reheating in detail. Finally, the accident mechanism could be applied to other realisations of inflation.

### v2

@@ 5 Topological defects
Magnetic monopoles can appear if the second homotopy group of the vacuum manifold is nonzero. When $G=\mathrm{SO}(3)$ is broken to $\mathrm{SO}(2)$ by a triplet VEV, as is the case in our model, this gives rise to the celebrated ’t Hooft-Polyakov monopole [43, 44] with $\pi_{2}(\mathrm{SO}(3)/\mathrm{SO}(2))=\pi_{2}[S^{2}]=\mathbb{Z}$. However, in our models, the relevant phase transition takes place before inflation, and, therefore, no monopoles are likely to remain in our Hubble patch. To be precise, symmetry breaking in our model proceeds in two steps: First, at the scale $v$, the symmetry group $G$ is broken completely by the 5-plet VEV, while the triplet VEV is zero in this phase. The vacuum manifold (or rather the scalar field manifold which minimizes the potential for all fields except the slowly-rolling inflaton) is $G$, and in different causally connected regions of the Universe, the scalar fields will take VEVs in different domains of $G$. However, the subsequent phase of inflation stretches out these inhomogeneities, such that, at all scales relevant to us, the $5$-plet VEV becomes effectively a single point in field space. Nontrivial field configurations which might evolve into monopoles are “inflated away”. The second step of symmetry breaking happens at the waterfall transition at the end of inflation, when the $5$-plet VEV relaxes to zero and a triplet VEV is generated, thereby restoring an $\mathrm{SO}(2)$ symmetry. The vacuum manifold becomes $\mathrm{SO}(3)/\mathrm{SO}(2)\simeq S^{2}$; however, no new monopoles can be created, since this would require energies sufficiently high to restore the entire $\mathrm{SO}(3)$ symmetry. We conclude that, for better or worse, magnetic monopoles will not affect the cosmology of our model.
@@
Cosmic strings can appear if the scalar vacuum manifold is not simply connected. This is not the case in the minimal model, but it is interesting to consider a non-minimal extension to study the phenomenological consequences. Consider a version of the model in which the triplet $\chi$ is a complex scalar field transforming under $G=\mathrm{SU}(2)\times\mathrm{U}(1)_{\chi}$, with $\mathrm{U}(1)_{\chi}$ acting as a phase rotation. All or part of $G$ may be thought of as gauged. We can further extend the model by promoting also $\phi$ to a complex field and gauging $\mathrm{U}(1)_{\phi}$, which is appealing because then the model no longer relies on any ad-hoc discrete symmetries, the role of the $\mathbb{Z}_{2}$ in the original model now being played by $\mathrm{U}(1)_{\phi}$. The technical details of these model versions are worked out in Appendix B. These models do allow for additional quartic scalar couplings, but the qualitative features of the inflaton potential, as well as CMB predictions, do not change: Crucially, there is still an accidentally light scalar field which can play the role of the inflaton, with a cosinusoidal loop-induced potential analogous to Eq. (6). Likewise, the effective mass parameter of the waterfall field $\chi$ depends on the value of the inflaton, and parameters can be chosen such that it becomes tachyonic at the end of inflation. The waterfall field is zero along the inflationary trajectory, so $\mathrm{U}(1)_{\chi}$ remains unbroken during inflation while $\mathrm{SU}(2)\times\mathrm{U}(1)_{\phi}$ is completely broken. The true vacuum is at $\langle\phi\rangle=0$ and $\langle|\chi|\rangle=\langle|\chi^{3}|\rangle=v_{\chi}\neq 0$, so the vacuum manifold is $S^{2}\times\mathrm{U}(1)_{\chi}$.
@@
where $v_{\chi}$ is the breaking scale of $\mathrm{U}(1)_{\chi}$. It is well known that this network quickly reaches a scaling regime, where the fraction of energy density stored in the CSs redshifts in the same way as the energy density of the background [46, 45]. Strings can intercommute and produce loops, which oscillate under their own tension and radiate energy in the form of GWs. Such a signal of GWs appears to us as a SGWB with a characteristic scale-invariant spectrum (if standard cosmology is assumed [47]). The peak and the amplitude of the GWs signal for stable strings are both given by the string tension, which, in order to match the signal recently reported by PTA collaborations, has to be $G_{\mathrm{\scriptscriptstyle{N}}}\,\mu\simeq 10^{-10}$ [15, 48], even though stable CSs seem to be disfavoured with respect to other cosmological sources [15]. The reason is that, for stable CSs, both the amplitude and the shape of the GW signal are controlled by $G_{\mathrm{\scriptscriptstyle{N}}}\mu$; once we fix the string tension to match the amplitude of the reported signal, the CS spectrum turns out to be flat in the nHz region, with data preferring a blue-tilted spectrum. $G_{\mathrm{\scriptscriptstyle{N}}}\mu\simeq 10^{-10}$ represents a trade-off between a signal which is not too weak and one which is not too flat. However, as the computation of the expected signal of GWs from a CS network is subject to large uncertainties, we can take as upper bound on the string tension the one given by the CMB constraint on the impact of CSs on the CMB power spectrum [49], $G_{\mathrm{\scriptscriptstyle{N}}}\mu\lesssim 10^{-7}$, which corresponds to an inflationary scale $V_{0}^{1/4}\lesssim\left(\lambda_{\chi}+\lambda_{\chi}^{\prime}\right)\left(3\times 10^{14}\right)\,{\rm GeV}$.
Finally, domain walls can appear if the scalar vacuum manifold is not connected. This, again, is not the case in our minimal model. In the following, we exploit a non-minimal model with disconnected minima, which is described in detail in Appendix C. In this case, the PTA signal can be fit with GWs produced from annihilating DWs, as explained in the next section. The waterfall field $\chi$ is taken to be a complex scalar triplet, subject to a discrete symmetry $\mathbb{Z}_{4\chi}:\chi\,\rightarrow\,i\chi$. This leads to disconnected minima which are degenerate, corresponding to stable DWs. However, the vacuum degeneracy can be easily lifted by softly breaking the $Z_{4\chi}$ discrete symmetry, allowing for the DWs to annihilate. Also in this model the inflationary predictions are the same as those for the minimal model.
@@ 5.1 Gravitational waves from unstable domain walls
In the model described in Appendix C, after the end of inflation the $\chi$ scalar potential has two disconnected, degenerate minima, so that DWs are formed with a surface energy density [45]
@@
However, the discrete symmetry ensuring the vacuum degeneracy may be softly broken, which lifts the degeneracy and introduces a potential bias $\Delta V$; see Eq. (C6). Provided that $\Delta V<0.795\,V_{0}$ [50], where $V_{0}$ is the potential energy at $\chi=0$, DWs are formed, but they are unstable as long as $\Delta V\neq 0$. The different domains annihilate when the bias pressure $\Delta V$ dominates over the DW surface tension at
@@
with $C_{d}\simeq 3$ [52], until all the observable Universe lies in the true vacuum. We require the DW network to disappear before the bias energy density of the false-vacuum patches dominates the Universe:
@@
Here $R\left(t_{\rm ann}\right)$ and $R\left(t_{0}\right)$ are the scale factors at the time of DW annihilation and today, respectively, $g_{*}$ is the number of effective degrees of freedom in relativistic energy, and $T_{\rm ann}$ is the temperature of the Universe at the moment of DW annihilation:
@@
Numerical simulations have shown that $b,c\simeq 1$ for $\mathbb{Z}_{2}$-type DWs [51], and our model falls in this category, while causality fixes $a=3$ [53]. [^7] In principle, the causality tail of the GW spectrum is affected by the QCD phase transition and the subsequent change in the number of relativistic degrees of freedom. This leads to a deviation from the pure power-law behaviour with $a=3$. Including the effect of the QCD phase transition is beyond the scope of this work, and we refer to Ref. [54] for a detailed discussion on the topic. By plugging Eqs. (36) and (39) into Eq. (38), we observe that the GW peak frequency scales with the DW network parameters as $f_{p}\sim\sqrt{\Delta V/\sigma}$. Similarly, by plugging Eq. (41) into Eq. (40), we find that $h^{2}\Omega_{\rm GW}(f_{p})\sim(\sigma^{2}/\Delta V)^{2}$.
By taking the quartic couplings to be of the order of one, the DW tension is related to the inflationary scale as $V_{0}\simeq\sigma^{4/3}$. This allows us to relate the GW signal generated by the DW network to inflation. It has been shown [15, 52, 55] that a network of unstable DWs is a good candidate to explain the signal of a SGWB recently detected by the NANOGrav Collaboration [15]. We find that a model with $V_{0}^{1/4}\simeq 2\times 10^{5}\,{\rm GeV}$ and a suitable bias $\Delta V$ leads to a GW signal which fits the NANOGrav detection. The peak of the GW signal is within the reach of the upcoming space-borne Laser Interferometer Space Antenna (LISA) [56] for $3\times 10^{7}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 2\times 10^{10}\,{\rm GeV}$, while for $3\times 10^{10}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 7\times 10^{12}\,{\rm GeV}$ the SGWB might be detected by the ET in the future. Fig. 5 shows several different benchmark GW spectra together with the sensitivity curves of current and planned experiments.
The energy density released by DW annihilation can potentially alter the expansion rate of the Universe. In order not to spoil BBN, the products of DW annihilation have to decay before the BBN epoque, either into dark radiation (DR) or into Standard Model (SM) particles [57]. Indeed, $\chi$ can decay into the $\mathrm{SO}(2)$ dark photon, or into SM degrees of freedom via, e.g., a Higgs-portal coupling. The dominant decay channel depends on the relative magnitude of the couplings and is related to the reheating mechanism, which we do not explore in this paper. The usual assumption, which also leads to the weakest constraints, is that such decays are already active at $T_{\rm ann}$ [57]. Assuming the DW energy density is transferred entirely into DR, the extra number of neutrinos species receives a correction [58]
@@ 6 Conclusions
Models with accidentally light scalars [3] are particularly well suited for realizing hybrid inflation. They feature tree-level flat directions, lifted by loop corrections, along which slow-roll inflation takes place. They also feature tree-level couplings of the inflaton to other scalar fields, which can become tachyonic along the inflationary direction, thereby ending inflation by the waterfall mechanism.
In this paper, we have presented a simple yet realistic example. It is defined by a symmetry group $G=\mathrm{SO}(3)$ and two scalar multiplets $\phi$ and $\chi$, transforming in the representation $\boldsymbol{5}$ and $\boldsymbol{3}$, respectively. The two fields may also carry an abelian charge (discrete or continuous), depending on the variant of the model. Gauging $\mathrm{SO}(3)$ and allowing for all renormalizable potential terms compatible with the symmetry, one degree of freedom $a$ in $\phi$ remains massless at tree level: This is what we call an accidentally light field, which we identify with the inflaton. The inflationary potential arises at the one-loop level, induced by the gauge coupling and a quartic coupling between $\phi$ and $\chi$. It has a form similar to that of Natural Inflation [4]. One component of $\chi$ plays the role of the waterfall field: Its mass-squared starts out positive and decreases during inflation until, at some critical point along the inflationary direction, it turns tachyonic. This destabilizes the waterfall field, which terminates inflation by fast-rolling down its steep potential.
The flatness of the inflationary potential is naturally protected: It is guaranteed by symmetry at the tree level, and it gets lifted by one-loop radiative corrections. Unlike pseudo-Goldstone inflatons, here the inflaton has tree-level non-derivative couplings to other scalars, which are necessary for an efficient waterfall. These couplings are typically a concern in earlier realizations of hybrid natural inflation, as they spoil the flatness, unless they are forbidden by some complicated additional symmetries or fine-tuned.
@@
Inflation ends as the waterfall field becomes tachyonic, and the Universe is reheated with a reheating temperature of the order of the inflationary scale, $T_{\rm RH}\sim{\cal O}(V_{0}^{1/4})$. A negative curvature of the potential triggers tachyonic instability of the waterfall field quantum fluctuations, which are exponentially enhanced. This leads to production of GWs, which reach us as a SGWB signal. The peak frequency of the GW spectrum is typically as large as $\mathcal{O}\left(10^{10}\right)\,{\rm Hz}$, but it can be lowered by sufficiently decreasing the inflationary scale. We find that the signal of GWs produced during tachyonic instability is within the reach of the Einstein Telescope for an inflationary scale in the range $3\times 10^{5}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 5\times 10^{11}{\rm GeV}$. Such a low inflationary scale, despite pointing to some UV completion well below $M_{\mathrm{P}}$, is perfectly viable and is associated to an unobservably small tensor-to-scalar ratio. This confirms that GWs from a tachyonic instability can act as a possible probe of low-scale inflationary scenarios.
The inflationary phase may be followed by a production of topological defects, depending on the specific symmetries which undergo SSB. In a model variation with $\phi$ and $\chi$ charged under an additional $U(1)$ gauge symmetry, the vacuum manifold is such that CSs are produced during the waterfall transition. The tension of the string network is bounded by CMB, and this constrains the inflationary scale to be $V_{0}^{1/4}\lesssim$ a few ×$\times 10^{14}\,{\rm GeV}$. Another model variation features a softly-broken $\mathbb{Z}_{4}$ symmetry acting on $\phi$ and $\chi$. In this case, the vacuum manifold is disconnected, and unstable DWs are generated at the end of inflation. They produce a SGWB, which can fit the signal recently detected by the NANOGrav Collaboration [15], albeit for an extremely low inflationary scale, $V_{0}^{1/4}\simeq 2\times 10^{5}\,{\rm GeV}$, assuming order one quartic couplings. The DW annihilation signal can also be relevant for future detection by LISA (ET) for $3\times 10^{7}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 2\times 10^{10}\,{\rm GeV}$ ($3\times 10^{10}\,{\rm GeV}\lesssim V_{0}^{1/4}\lesssim 7\times 10^{12}\,{\rm GeV}$).
Upcoming experiments will soon be able to further constrain or rule out our models, for example, by measuring the sign of $n_{r}$, or by a discovery of a SGWB by the LIGO-Virgo-KAGRA network. Concerning future directions, it would be interesting to couple our models to the Standard Model in order to study reheating in detail. Finally, the accident mechanism could be applied to other realizations of inflation.